a committee of 3 people is to be selected from four women and five men the rules state that there must be at least one man and one women on the committee in how many ways can the committee be chosen
Answer by edjones(8007)

1 man 2 women: nCr=5C1=5, 4C2=6; 6*5=30
2 men 1 woman; 5C2=10, 4C1=4; 10*4=40
40+30=70 ways
.
Ed